OpenStudy (anonymous):

Which of the following is a solution of x2 - 8x = -27?

OpenStudy (anonymous):

negative 4 minus i square root of 11 4 plus i square root of 11 negative 4 minus i square root of 44 4 plus i square root of 44

OpenStudy (vishweshshrimali5):

\[\large{-4 - i\sqrt{11}}\] \[\large{4+i\sqrt{11}}\] \[\large{-4-i\sqrt{44}}\] \[\large{4+i\sqrt{44}}\]

OpenStudy (vishweshshrimali5):

These are the options ?

OpenStudy (anonymous):

correct

OpenStudy (vishweshshrimali5):

Okay what do you think the answer should be and why ?

OpenStudy (vishweshshrimali5):

Have you tried it ?

OpenStudy (anonymous):

I have but I think im trying it wrong

OpenStudy (vishweshshrimali5):

The equation given in the question is: \[\large{x^2 - 8x = -27}\] \[\large{\implies x^2 - 8x + 27 = 0}\] Okay what did you try ?

OpenStudy (vishweshshrimali5):

You can compare this equation with the general quadratic equation: \[\large{ax^2 + bx + c = 0}\] On comparing, you can find out a,b and c. Using these values and using quadratic formula we can find out the solution of the equation: \[\large{x = \cfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}}\]

OpenStudy (anonymous):

ohh I wrote the number different hold on let me try that

OpenStudy (vishweshshrimali5):

Sure :)

OpenStudy (anonymous):

so I tried but im still clueless

OpenStudy (vishweshshrimali5):

Its okay. What have you tried ?

OpenStudy (anonymous):

what I tried was wrong, how should I solve this?

OpenStudy (vishweshshrimali5):

Well follow this method: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 You can compare this equation with the general quadratic equation: \[\large{ax^2 + bx + c = 0}\] On comparing, you can find out a,b and c. Using these values and using quadratic formula we can find out the solution of the equation: \[\large{x = \cfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}}\] \(\color{blue}{\text{End of Quote}}\) Step by step First find out a,b and c

OpenStudy (anonymous):

how do I find that? can u write the equation and I solve it?

OpenStudy (vishweshshrimali5):

Okay lets see: \[\large{x^2 - 8x + 27 = 0}\] \[\large{ax^2 + bx + c = 0}\]

OpenStudy (vishweshshrimali5):

Now lets compare both equations

OpenStudy (anonymous):

alright

OpenStudy (vishweshshrimali5):

In first equation coefficient of x^2 = 1 In second it is = a So, a = 1

OpenStudy (vishweshshrimali5):

Now similarly find out b

OpenStudy (anonymous):

b is 8? c is 27?

OpenStudy (vishweshshrimali5):

c is correct Check the sign of b again

OpenStudy (vishweshshrimali5):

In first equation coefficient of x = -8

OpenStudy (vishweshshrimali5):

Don't forget the minus sign

OpenStudy (anonymous):

okay so b is -8

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (anonymous):

You now have the a , b , and c to use in your quadratic formula.

OpenStudy (anonymous):

@ChristopherToni

OpenStudy (anonymous):

can you help me? @ChristopherToni

OpenStudy (anonymous):

Up to this point, you've determined that a=1, b=-8 and c=27. Plugging this into the quadratic formula gives us \(\large x=\dfrac{-(-8) \pm \sqrt{(-8)^2-4(1)(27)}}{2(1)}\). What do you get when you simplify this? :-)

OpenStudy (anonymous):

I troed and I got